# Plasma acceleration of polarized particle beams

### v1

@@ 5.3 Application: deciphering in situ electron dynamics of ultrarelativistic plasma
Figure 13: The schematic for $\gamma$-ray photon emissions from a plasma interacting with a laser pulse. (b) The LP orientation is along the local azimuthal direction. The analytically predicted LP orientation for the electron undergoing acceleration (c) and deceleration (d). The simulated $\gamma$-photon LP for the time at (e) $t=10$ and (f) 40 fs, with the LP degree $P_{\mathrm{LP}}$ and the normalized number distribution $d^{2}N_{ph}/\sin\theta d\theta d\phi$. (g) The time evolution of the electron energy spectrum $dN_{e}/d\varepsilon_{e}$, the polarization orientation $\delta\phi$, and the kinetic energy $\varepsilon_{e}$. Reproduced under the terms of the CC-BY license from [82]. Copyright 2022, The Authors, published by American Physical Society.
In the moderate QED regime $\chi_{e}<1$, the direction of the emitted $\gamma$-photon LP is primarily parallel to the acceleration direction perpendicular to the electron momentum, $\textbf{a}_{\perp}\equiv\textbf{a}-(\textbf{a}\cdot\hat{\textbf{v}})\hat{\textbf{v}}$, where the hat symbol denotes the unit vector. Thus, the polarization orientation is derived as $\hat{a}_{\perp,y}\approx-\sin\phi\left(\frac{\Gamma}{\epsilon\gamma_{e}}+\frac{\kappa_{b}\cos\theta}{\epsilon\Gamma}\right)-\frac{(-\bm{\beta}\cdot\textbf{E})\cos\phi}{\gamma_{e}}$ and $\hat{a}_{\perp,z}\approx\cos\phi\left(\frac{\epsilon\Gamma}{\gamma_{e}}+\frac{\epsilon\kappa_{b}\cos\theta}{\Gamma}\right)-\frac{(-\bm{\beta}\cdot\textbf{E})\sin\phi}{\gamma_{e}}$, where $\bm{\beta}=\textbf{v}/c$, $\theta=\arctan[(p_{y}^{2}+p_{z}^{2})^{1/2}/p_{x}]$ and $\phi=\mathrm{arctan2}(p_{z},p_{y})$. When the electron energy gain is negligible, i.e. $-\bm{\beta}\cdot\textbf{E}=0$, the orientation of the $\gamma$-photon LP is along the azimuthal direction $\hat{\textbf{a}}_{a}=(-\sin\phi,\cos\phi)^{\intercal}$, which collectively resembles multiple concentric rings with each polarization segment along the azimuthal direction [see Fig. 13(b)]. The deviation of the $\gamma$-photon LP orientation from the azimuthal direction is quantified by $\delta\phi\in[-90^{\circ},90^{\circ}]$, which is the relative angle between $\hat{\textbf{a}}_{\perp}$ and $\hat{\textbf{a}}_{a}$ and is calculated as
@@
If the radiating electron is undergoing acceleration with $-\bm{\beta}\cdot\textbf{E}>0$ ($-\bm{\beta}\cdot\textbf{E}<0$), the GPP orientation $\delta\phi>0$ ($\delta\phi<0$) corresponds to the counter-clockwise (clockwise) spiral tendency in the angular distribution of $\gamma$-photon LP as shown in Fig. 13(c) [Fig. 13(d)]. The 3D-PIC simulation results of the orientation of the emitted $\gamma$-photon LP exhibit a counter-clockwise spiral tendency at $t=10$ fs [Fig. 13(e)] reproducing well the analytical prediction for the accelerating electron. Here, the averaged polarization angle is $\delta\phi\approx 30.6^{\circ}$ and the LP degree $P_{\mathrm{LP}}\approx 15.3\%$. The clockwise spiral tendency in Fig. 13(f) implies the deceleration of plasma electrons occurring later at $t=40\,$fs. The time-resolved $\delta\phi$ explicitly reflects the electrons being predominantly accelerated (decelerated) at $t\lesssim t_{s}^{\varepsilon_{e}}$ ($t\gtrsim t_{s}^{\varepsilon_{e}}$) [see Fig. 13(g)], where $t_{s}^{\varepsilon_{e}}\sim 35\,$fs is the electron energy saturation time. The correlation between the spiral tendency of photon LP $\delta\phi$ and electron in-situ dynamics demonstrates that the retrieving method of utilizing the information of photon polarization might be beneficial for better understanding of phenomena in broad high-intensity interaction scenarios including ion acceleration, direct laser acceleration, high-harmonic generation, brilliant photon emission, ultra-dense nanopinches, and $e^{-}e^{+}$ pair plasma cascades.
5.4 Application: exploring electron slingshot acceleration in relativistic preturbulent shocks
@@
As a versatile information carrier of multi-messenger astrophysics [101, 102, 103], photon polarization is critical for measuring the magnetic configuration nearby black holes [104] and crab nebulae [105], and for analyzing the particle acceleration in the blazar’s jet [106]. Recent studies about relativistic collisionless shocks (RCSs) have attracted great attention since RCSs could instigate electron stochastic acceleration, akin to the Fermi process [107, 108], which has been well recognized as sources of energetic electrons in the universe. Therefore, the question arises whether the polarization feature of spontaneously emitted photons can be employed to reveal some new acceleration mechanisms in a RCS.
@@
Figure 14: The dynamics of a counter-streaming plasma interface: The electron density $n_{e}$ at $t=80\pi/\omega_{pe}$ for the case with ion fraction of (a) $\eta=0.4$ and (b) $\eta=0.01$, where lines present the typical electron moving tendency with stars marking the photon emission and the histograms display the spatial distribution of emitted photons with $\omega_{ph}>10^{-2}\omega_{ph}^{m}$. (c) Photon LP degree $\mathcal{Q}\right>$ and energy spectra $\omega_{ph}dN_{ph}/d\omega_{ph}$ vs $\omega_{ph}$. (d) $\mathcal{Q}\right>$ and $dN_{ph}/d\theta_{ph}$ vs $\theta_{ph}$. Reproduced under the terms of the CC-BY license from [109]. Copyright 2023, The Authors, published by American Physical Society.
@@
In a latest study, Gong et al. [109] found that photon polarization information can be utilized to explore the new mechanism of electron slingshot acceleration in counter-streaming plasma. They focused on the transient electron dynamics in the transition to turbulence nearby the counter-streaming interface of an unmagnetized pair-loaded relativistic collisionless shock precursor, which is potentially associated with the outflow of GRBs.
@@
Gong et al. [109] carried out 2D simulations of counter-streaming RCSs, see Fig. 14. The snapshot of the electron density $n_{e}$ in Fig. 14(a) exhibits that the filamentation exclusively exists at the front of the RCS interface. Between two adjacent filaments, an electron focusing point emerges, and following that, two oblique density strips stretch out [see Fig. 14(b)]. Behind the strips, the coherent filaments and focusing points disappear while the turbulence shows up. It has to be highlighted that the photons with energy $\varepsilon_{ph}\equiv\hbar\omega_{ph}>10^{-2}\hbar\omega_{ph}^{m}$ are primarily emitted by electrons nearby the interface, where $\omega_{ph}^{m}\sim 10^{8}\omega_{pe}$ is the photon cut-off frequency and $\hbar$ Planck’s constant. The degree of photon’s linear polarization along the direction of the electron’s transverse acceleration is characterized by the Stokes parameter $\mathcal{Q}$ [110], formulated as
@@
where $K_{n}(\zeta)$ is the modified secondary Bessel function, $\widetilde{K}_{1/3}(\zeta)=\int_{\zeta}^{\infty}K_{1/3}(z)\mathrm{d}z$, $\zeta=2\varepsilon_{ph}/[3\chi_{e}(\varepsilon_{e}-\varepsilon_{ph})]$, and $\varepsilon_{e}=\gamma_{e}m_{e}c^{2}$ the electron energy; $\chi_{e}\equiv(e\hbar/m_{e}^{3}c^{4})|F_{\mu\nu}p^{\nu}|$ is the electron quantum strong-field parameter with the field tensor $F_{\mu\nu}$, and the electron four-momentum $p^{\nu}$. At $\chi_{e}\ll 0.1$, $\partial\mathcal{Q}/\partial\varepsilon_{ph}>0$ predicted by Eq. (13) manifests a monotonic dependence of $\mathcal{Q}$ on $\omega_{ph}$, because for the higher-frequency radiation the formation length is shorter and the preservation of the local polarization degree is improved. This monotonic dependence is confirmed by the results of $\eta=0.01$ [see Fig. 14(c)(d)], where electrons experience stochastic acceleration and the photon emission is isotropic in the angular space. However, for $\eta=0.4$ [see Fig. 14(c)(d)], the averaged polarization degree $\mathcal{Q}\right>$ versus $\omega_{ph}$ exhibits non-monotonic dependence, with a polarization dip $\Delta\mathcal{Q}_{\omega}\approx 4.5\%$ and a bandwidth ratio $\mathcal{B}_{\omega}\equiv\omega_{ph}^{m}/\omega_{ph}^{*}\sim 10^{3}$, contradictory to the aforementioned monotonic dependence. Here, $\omega_{ph}^{*}$ is the local maximum point of the function $\mathcal{Q}\right>$ vs $\omega_{ph}$ [see Fig. 14(c)]. Utilizing the correlation between the work contribution and the non-monotonic dependence of photon polarization, Gong et al. [109] identified the mechanism of electron slingshot acceleration.
@@ 5.5 Vortex gamma-ray photons
Vortex gamma-ray photons are high-energy photons that carry intrinsic orbital angular momentum (OAM) in addition to their intrinsic spin angular momentum [111, 112, 113, 114, 115]. Characterized by a helical wavefront structure, these photons exhibit a quantized twist, allowing them to have unique properties that differ from ordinary gamma rays. Vortex photons have been successfully generated across the visible to X-ray range using techniques such as optical mode conversion, high harmonic generation, and coherent radiation from helical undulators or laser facilities [116, 117, 118, 119]. As the new degree of freedom of information is carried by the vortex state, vortex gamma-ray photons have broad applications in various fields, including astrophysics, strong-laser physics, particle physics, and nuclear physics [120, 121, 122, 123, 124]. However, the generation of vortex gamma photons remains a big challenge.
With the rapid progress of ultraintense laser techniques, the theoretical schemes of generating vortex gamma-ray photons have been extensively proposed, where most of the studies utilize Compton backscattering to obtain high energy and OAM gamma-ray photons [125, 111, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135]. Nevertheless, most of these studies have only considered vortex beams with collective OAM in laser-plasma interactions, the intrinsic OAM carried by gamma-ray photons is still unclear.
@@
Recently, a quantum electrodynamical scattering theory has been considered to characterize the multiphoton absorption and the process of angular momentum transfer in the generation of vortex gamma-ray photons. For instance, Ababekri et al. [136] explored the generation of vortex gamma photons through nonlinear Compton scattering of ultrarelativistic electrons in a circularly polarized laser pulse. They discuss the vortex phase structure of the scattering matrix element, the transformation of the vortex phase to the emitted photon, and the radiation rate for vortex gamma photons. Guo et al. [137] investigated the production of gamma-ray photons with large intrinsic OAM via nonlinear Compton scattering. Gamma-ray photons carrying controllable OAM quantum numbers from tens to thousands of units are generated when employing vortex electrons at moderate laser intensities. Jiang et al. [138] studied the generation of vortex gamma photons with controllable spin and orbital angular momenta via nonlinear Compton scattering of two-color counter-rotating CP laser fields. They find that the polarization and vortex charge of the emitted gamma photons can be controlled by tuning the relative intensity ratio of the two-color CP laser fields. Utilizing the quantum radiation theory could lead to a better understanding of the properties of the emitted vortex gamma-ray photons, which potentially opens new opportunities in hadron, atomic, nuclear, particle, and high-energy physics.

### v2

@@ 5.3 Application: deciphering in situ electron dynamics of high-intensity laser-driven plasma
Figure 13: The schematic for $\gamma$-ray photon emissions from a plasma interacting with a laser pulse. (b) The LP orientation is along the local azimuthal direction. The analytically predicted LP orientation for the electron undergoing acceleration (c) and deceleration (d). The simulated $\gamma$-photon LP for the time at (e) $t=10$ and (f) 40 fs, with the LP degree $P_{\mathrm{LP}}$ and the normalized number distribution $d^{2}N_{ph}/\sin\theta d\theta d\phi$. (g) The time evolution of the electron energy spectrum $dN_{e}/d\varepsilon_{e}$, the polarization orientation $\delta\phi$, and the kinetic energy $\varepsilon_{e}$. Reproduced under the terms of the $C$-BY license from [111]. Copyright 2022, The Authors, published by American Physical Society.
In the moderate QED regime $\chi_{e}<1$, the direction of the emitted $\gamma$-photon LP is primarily parallel to the acceleration direction perpendicular to the electron momentum, $\textbf{a}_{\perp}\equiv\textbf{a}-(\textbf{a}\cdot\hat{\textbf{v}})\hat{\textbf{v}}$, where the hat symbol denotes the unit vector. Thus, the polarization orientation is derived as $\hat{a}_{\perp,y}\approx-\sin\phi\left(\frac{\Gamma}{\epsilon\gamma_{e}}+\frac{\kappa_{b}\cos\theta}{\epsilon\Gamma}\right)-\frac{(-\bm{\beta}\cdot\textbf{E})\cos\phi}{\gamma_{e}}$ and $\hat{a}_{\perp,z}\approx\cos\phi\left(\frac{\epsilon\Gamma}{\gamma_{e}}+\frac{\epsilon\kappa_{b}\cos\theta}{\Gamma}\right)-\frac{(-\bm{\beta}\cdot\textbf{E})\sin\phi}{\gamma_{e}}$, where $\bm{\beta}=\textbf{v}/c$, $\theta=\arctan[(p_{y}^{2}+p_{z}^{2})^{1/2}/p_{x}]$ and $\phi=\mathrm{arctan2}(p_{z},p_{y})$. Here, $\chi_{e}=(e\hbar/m_{e}^{3}c^{4})|F_{\mu\nu}p^{\nu}|$ denotes the invariant quantum parameter of the electron with the field tensor $F_{\mu\nu}$ and the electron four-momentum $p^{\nu}$ and $\Gamma\equiv\gamma_{e}-(p_{x}/m_{e}v_{\mathrm{vph}})$ is the dephasing value of the electron within the laser field. $\kappa_{b}$ is the coefficient for characterizing the self-generated magnetic field and $\epsilon$ is the ellipticity of the laser electromagnetic wave. $p_{x,y,z}$ denotes the electron momentum.
When the electron energy gain is negligible, that is $\bm{\beta}\cdot\textbf{E}=0$, the orientation of the $\gamma$-photon LP is along the azimuthal direction $\hat{\textbf{a}}_{a}=(-\sin\phi,\cos\phi)^{\intercal}$, which collectively resembles multiple concentric rings with each polarization segment along the azimuthal direction [see Fig. 13(b)]. The deviation of the $\gamma$-photon LP orientation from the azimuthal direction is quantified by $\delta\phi\in[-90^{\circ},90^{\circ}]$, which is the relative angle between $\hat{\textbf{a}}_{\perp}$ and $\hat{\textbf{a}}_{a}$ and is calculated as
@@
If the radiating electron is undergoing acceleration with $\bm{\beta}\cdot\textbf{E}>0$ ($\bm{\beta}\cdot\textbf{E}<0$), the GPP orientation $\delta\phi>0$ ($\delta\phi<0$) corresponds to the counter-clockwise (clockwise) spiral tendency in the angular distribution of the $\gamma$-photon LP as shown in Fig. 13(c) [Fig. 13(d)]. The 3D-PIC simulation results of the orientation of the emitted $\gamma$-photon LP exhibit a counter-clockwise spiral tendency at $t=10$ fs [Fig. 13(e)] reproducing well the analytical prediction for the accelerating electron. Here, the averaged polarization angle is $\delta\phi\approx 30.6^{\circ}$ and the LP degree is $P_{\mathrm{LP}}\approx 15.3\%$. The clockwise spiral tendency in Fig. 13(f) implies the deceleration of plasma electrons occurring later at t=40t=40\,fs. The time-resolved $\delta\phi$ explicitly reflects that the electrons are predominantly accelerated (decelerated) at $t\lesssim t_{s}^{\varepsilon_{e}}$ ($t\gtrsim t_{s}^{\varepsilon_{e}}$) [see Fig. 13(g)], where $t_{s}^{\varepsilon_{e}}\sim 35\,$fs is the saturation time of the electron energy. The correlation between the spiral tendency of photon LP $\delta\phi$ and electron in-situ dynamics demonstrates that the retrieving method of utilizing the information of photon polarization might be beneficial for better understanding of phenomena in broad high-intensity interaction scenarios including ion acceleration, direct laser acceleration, high-harmonic generation, brilliant photon emission, ultra-dense nanopinches, and $e^{-}e^{+}$ pair plasma cascades.
5.4 Application: exploring electron dynamics in relativistic astrophysical plasmas
@@
As a versatile information carrier of multi-messenger astrophysics [134, 135, 136], photon polarization is critical to measuring the magnetic configuration near black holes [137] and crab nebulae [138], and for analyzing the particle acceleration in the blazar jet [139]. Recent studies on relativistic collisionless shocks (RCSs) have attracted great attention since RCSs could instigate electron stochastic acceleration, similar to the Fermi process [140, 141], which has been well recognized as sources of energetic electrons in the universe. Therefore, the question arises of whether the polarization feature of spontaneously emitted photons can be employed to reveal some new acceleration mechanisms in an RCS.
@@
Figure 14: The dynamics of a counter-streaming plasma interface: The electron density $n_{e}$ at $t=80\pi/\omega_{pe}$ for the case with ion fraction of (a) $\eta=0.4$ and (b) $\eta=0.01$, where lines present the typical electron moving tendency with stars marking the photon emission and the histograms display the spatial distribution of emitted photons with $\omega_{ph}>10^{-2}\omega_{ph}^{m}$. (c) Photon LP degree $\left<\mathcal{Q}\right$> and energy spectra $\omega_{ph}dN_{ph}/d\omega_{ph}$ vs $\omega_{ph}$. (d) $\left<\mathcal{Q}\right$> and $dN_{ph}/d\theta_{ph}$ vs $\theta_{ph}$. Reproduced under the terms of the $C$-BY license from [142]. Copyright 2023, The Authors, published by American Physical Society.
@@
In a recent study, Gong et al. [142] found that photon polarization information can be utilized to explore the new mechanism of electron slingshot acceleration in counter-streaming plasma. They focused on the transient electron dynamics in the transition to turbulence near the counter-streaming interface of an unmagnetized pair-loaded relativistic collisionless shock precursor, which is potentially associated with the outflow of gamma-ray bursts (GRBs).
@@
Gong et al. [142] carried out 2D simulations of counter-streaming RCSs, see Fig. 14. The snapshot of the electron density $n_{e}$ in Fig. 14(a) shows that the filamentation exists at the front of the RCS interface. Between two adjacent filaments, an electron focusing point emerges, and following that, two oblique density strips stretch out [see Fig. 14(b)]. Behind the strips, the coherent filaments and focusing points disappear, while the turbulence shows up. It should be highlighted that photons with energy $\varepsilon_{ph}\equiv\hbar\omega_{ph}>10^{-2}\hbar\omega_{ph}^{m}$ are primarily emitted by electrons near the interface, where $\omega_{ph}^{m}\sim 10^{8}\omega_{pe}$ is the photon cut-off frequency and $\hbar$ is Planck’s constant. The degree of photon’s linear polarization along the direction of the electron’s transverse acceleration is characterized by the Stokes parameter $\mathcal{Q}$ [143], formulated as
@@
where $K_{n}(\zeta)$ is the modified secondary Bessel function, $\widetilde{K}_{1/3}(\zeta)=\int_{\zeta}^{\infty}K_{1/3}(z)\mathrm{d}z$, $\zeta=2\varepsilon_{ph}/[3\chi_{e}(\varepsilon_{e}-\varepsilon_{ph})]$, and $\varepsilon_{e}=\gamma_{e}m_{e}c^{2}$ the electron energy; $\chi_{e}\equiv(e\hbar/m_{e}^{3}c^{4})|F_{\mu\nu}p^{\nu}|$ is the electron quantum strong-field parameter with the field tensor $F_{\mu\nu}$, and the electron four-momentum $p^{\nu}$. At $\chi_{e}\ll 0.1$, $\partial\mathcal{Q}/\partial\varepsilon_{ph}>0$ as predicted by Eq. (13) manifests a monotonic dependence of $\mathcal{Q}$ on $\omega_{ph}$, because for the higher-frequency radiation the formation length is shorter and the preservation of the local polarization degree is improved. This monotonic dependence is confirmed by the results of $\eta=0.01$ [see Fig. 14(c)(d)], where electrons experience stochastic acceleration and photon emission is isotropic in the angular space. However, for $\eta=0.4$ [see Fig. 14(c)(d)], the averaged polarization degree $\left<\mathcal{Q}\right$> versus $\omega_{ph}$ exhibits non-monotonic dependence, with a polarization dip $\Delta\mathcal{Q}_{\omega}\approx 4.5\%$ and a bandwidth ratio $\mathcal{B}_{\omega}\equiv\omega_{ph}^{m}/\omega_{ph}^{*}\sim 10^{3}$, contradictory to the aforementioned monotonic dependence. Here, $\omega_{ph}^{*}$ is the local maximum point of the function $\left<\mathcal{Q}\right$> vs $\omega_{ph}$ [see Fig. 14(c)]. Using the correlation between the work contribution and the non-monotonic dependence of photon polarization, Gong et al. [142] identified the mechanism of electron slingshot acceleration.
Another recent study by Gong et al. [144] reported a understanding of anomalous gamma-ray photon polarization in radiation-dominated magnetic reconnection. Using particle-in-cell simulations that account for radiation reaction, spin dynamics, and quantum polarization effects, Gong et al. found the formation of spin-polarized condensed plasmoids. Electrons in the reconnection layer contract in phase space toward a spiral attractor induced by radiation damping, leading to the emergence of compact, high-density plasmoids with spin alignment along the local magnetic field. Compared with conventional plasmoids, these structures demonstrate much stronger density compression. More importantly, their radiation displays an unexpected linear polarization: the emitted gamma-ray photons are polarized perpendicular to the electron motion plane, in sharp contrast to classical synchrotron predictions. The study identifies spin-flip transitions as the key mechanism that reshapes angular momentum transfer and produces this anomalous polarization. This finding provides not only theoretical evidence for condensed spin-ordered plasmoids in magnetic reconnection, but also a new perspective for explaining puzzling polarization signals detected from extreme astrophysical environments such as magnetars [145, 146, 147].
@@ 5.5 Vortex gamma-ray photons
Vortex gamma-ray photons are high-energy photons that carry intrinsic orbital angular momentum (OAM) in addition to their intrinsic spin angular momentum [148, 149, 150, 151, 152, 153]. Characterized by a helical wavefront structure, these photons exhibit a quantized twist, allowing them to have unique properties that differ from ordinary gamma rays. Vortex photons have been successfully generated across the visible to X-ray range using techniques such as optical mode conversion, high harmonic generation, and coherent radiation from helical undulators or laser facilities [154, 155, 156, 157, 158]. As the new degree of freedom of information is carried by the vortex state, vortex gamma-ray photons have broad applications in various fields, including astrophysics, strong-field laser physics, particle physics, and nuclear physics [159, 160, 161, 162, 163]. However, the generation of vortex gamma photons remains a big challenge.
With the rapid progress of ultraintense laser techniques, most studies utilize Compton backscattering to obtain high-energy and OAM gamma-ray photons [164, 148, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174]. However, the produced OAM-dominated beams can be described semiclassically as ensembles of particles moving at an angle that carry classical OAM about the beam axis; a rotating cloud whose large mechanical OAM arises purely as a collective classical effect. In contrast, OAM in vortex states is an intrinsic property of individual particles, not a collective effect. From the classical electro-magnetic field radiation of energetic electron beams colliding with the vortex laser field, it is found that the accumulative field structure shows vortex features [175]. Inherent OAMs carried by gamma-photons have been shown in linear Compton scattering between energetic electrons and vortex laser photons, in the framework of quantum electrodynamics (QED) [176].
@@
In the nonlinear regime, the QED scattering theory including the quantum number of angular momentum has been considered to characterize multiphoton absorption and the process of angular momentum transfer in the generation of vortex gamma-ray photons. For instance, Bu et al. [153] showed that in Compton scattering of ultrarelativistic electrons in a CP laser pulse, absorption of spin-angular-momenta of multiple CP laser photons leads to efficient generation of twisted gamma photons in the weak linear regime and also OAM electrons in the strong linear case. Ababekri et al. [177] explored the generation of vortex gamma photons in similar setups. They discuss the vortex phase structure of the scattering matrix element, the transformation of the vortex phase into the emitted photon, and the radiation rate for vortex gamma photons. Guo et al. [178] investigated the production of gamma-ray photons with large intrinsic OAM via nonlinear Compton scattering. Gamma-ray photons carrying controllable OAM quantum numbers from tens to thousands of units are generated when employing vortex electrons at moderate laser intensities. Jiang et al. [179] studied the generation of vortex gamma photons with controllable spin and orbital angular momenta via nonlinear Compton scattering of two-color counter-rotating CP laser fields. They find that the polarization and vortex charge of the emitted gamma photons can be controlled by tuning the relative intensity ratio of the two-color CP laser fields. Alternatively, vortex gamma photons can also be efficiently generated through Bremsstrahlung, whether from initially twisted electrons [180] or through spin-to-orbital coupling with spin-polarized electrons [181].
Utilizing quantum radiation theory could lead to a better understanding of the properties of the emitted vortex gamma-ray photons, which potentially opens new opportunities in hadron, atomic, nuclear, particle, and high-energy physics. Further detailed discussion of the promises and challenges of high-energy vortex photons and other vortex particles can also be found in relevant review papers, e.g. Bliokh et al. [182], Lloyd et al. [183], Larocque et al. [184], and Ivanov [151].